how do you convert cm^2 and cm into dL?
the problem is (144.76g/cm^2)/ 0.3765cm = ? mg/dL

Answer by Alan3354(69443)

how do you convert cm^2 and cm into dL?
----------------
cm^2 is area, cm is distance. No conversion is possible.
dL is Liters/10, = 100 cm^3, which is volume. No conversion is possible.
-----
the problem is (144.76g/cm^2)/ 0.3765cm = ? mg/dL
===============
Pretending it makes sense:
(144.76g/cm^2)/ 0.3765cm = ? mg/dL
-----
1 dL = 100 cm^3
----
(144.76g/ 0.3765cm^3) = ? mg/dL
= 144760 mg/0.003765
=~ 38448871 mg/dL
-----------------